OpenStudy (anonymous):

Can someone help me with, 5(x + 6) 3x - 4 (Without giveing the answer)

Parth (parthkohli):

Is that an equals sign?

OpenStudy (anonymous):

<=

OpenStudy (anonymous):

srry forgot to fix it

Parth (parthkohli):

\[5(x + 6) \le 3x - 4 \]You start by distributing 5 into x + 6.

Parth (parthkohli):

Actually, you solve it exactly like you solve \(5(x + 6) = 3x - 4\)

OpenStudy (anonymous):

ok so do 5 * x and 5 * 6 and get 5x * 30

OpenStudy (anonymous):

and do -3x to both sides and get 5x * 30 = -4

OpenStudy (anonymous):

i mean 2x * 30 = -4

OpenStudy (anonymous):

That's a +, not a *

OpenStudy (anonymous):

so 2x +30??

OpenStudy (anonymous):

2x + 30 = -4. Then +4 to both sides then divide 2 to both sides? and get, x >= 15?

OpenStudy (anonymous):

listen the equation is 5(x + 6) 3x - 4 this onlyy????

OpenStudy (anonymous):

5(x + 6) <= x - 4

OpenStudy (anonymous):

No, you have to subtract 30 from both sides. If you add 4 to both sides you get 2x+34=0, you need the constant on the other side of the = from the variable.

OpenStudy (anonymous):

so 2x + 30 = -4. oh :PP -30 to both sides and divide 2 to both sides?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

giving me 2x = -34 divide 2 to both sides and get x >= -17?

OpenStudy (anonymous):

>= or <=?

OpenStudy (anonymous):

>= cause i have to flip the signs

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

or do i?